OpenStudy (anonymous):

Tried this one a couple times, I'll submit the answers I have, but my homework says I'm not getting ALL of them right Let f(x) = x+1 and g(x) = 1/{x+1}. Then (f of f) (x) = x^2+3+1 (f of g) (x) = x/(x^2+1) (g of f) (x) = 1/(x^3+x+1) (g of g) (x) = 1/(x^3+x^2)+1

OpenStudy (mathstudent55):

Start with the first one, f of f. There is no square in function f. How do you get an x^2 in the answer?

OpenStudy (anonymous):

I did x(x)+1, which got me x^2+1 for the first f. Am I doing that wrong?

jimthompson5910 (jim_thompson5910):

Start with f(x) = x+1

jimthompson5910 (jim_thompson5910):

everywhere you see an x, replace it with f(x) f(x) = x+1 f(x) = (x)+1 f( f(x) ) = ( f(x) )+1 then on the right side, replace f(x) with x+1, since f(x) = x+1 f( f(x) ) = ( f(x) )+1 f( f(x) ) = ( x+1 )+1 f( f(x) ) = x+1 +1 f( f(x) ) = x+2 So (f o f)(x) = x+2

OpenStudy (anonymous):

So you don't multiply them together, you just replace the x with the x?

jimthompson5910 (jim_thompson5910):

you replace x with f(x)

jimthompson5910 (jim_thompson5910):

let's use f(x) and g(x)

jimthompson5910 (jim_thompson5910):

(f o g) (x) is really f( g(x) )

jimthompson5910 (jim_thompson5910):

to calculate f( g(x) ), we start with f(x)